How would I solve this equation? I just really need the steps. Please give a walk through process.
4x + 7 = 25 - 5x
and
t + 9 = 3t - 5

Answer by stanbon(75887)

4x + 7 = 25 - 5x
Add 5x to both sides to get:
9x + 7 = 25
Subtract 7 from both sides to get:
9x = 18
Divide both sides by 9 to get:
x = 2
-----------------------------------------
and
t + 9 = 3t - 5
Subtract t from both sides to get:
9 = 2t -5
Add 5 to both sides
2t = 14
Divide both sides by 2 to get:
t = 7
=================
Cheers,
Stan H.